OpenStudy (anonymous):

Solve the system by any method. Please show your steps. y=x+2 y=x^2

OpenStudy (anonymous):

Substitute the expression for y from the first equation into the second equation and solve for x. Just as before.

OpenStudy (anonymous):

y=x^2+2?

OpenStudy (anonymous):

Not quite. From the first equation, we know that y=x+2. So in the second equation, put x+2 in place of y.

OpenStudy (anonymous):

The important point here is to recognize the meaning of the equals sign(=). The equation y=x+2 means that y and x+2 are exactly the same. That means that wherever we have y, we could replace it with x+2 and vice versa. Make sense?

OpenStudy (anonymous):

So the second equation can be written\[x+2=x ^{2}\]Can you solve this for x?

OpenStudy (anonymous):

yesfff

OpenStudy (anonymous):

Great. What do you get?

OpenStudy (anonymous):

2,-1

OpenStudy (anonymous):

Perfect. Now, take of these x-values and substitute them into either of the given equations to solve for y. What do you get?

OpenStudy (anonymous):

im not sure what the soultion would be? @ospreytriple

OpenStudy (anonymous):

Sorry, was away for a sec. Take the first equation, y=x+2. Evaluate it for x=2 and for x=-1. There will be two solutions. What are they?

OpenStudy (anonymous):

y=x*-1+2 ?

OpenStudy (anonymous):

i dont think that correct set up

OpenStudy (anonymous):

There is no need to use the variable x any more. You've already determined the value of x. Remember the meaning of the equals sign. By saying x=2, that means that we can replace x with 2. So the first equation becomes y=2+2 for x=2 and it becomes y=-1+2 for x=-1. Now, what do you get for y?

OpenStudy (anonymous):

y=1

OpenStudy (anonymous):

That's one solution. What's the other one? (when x=2)

OpenStudy (anonymous):

and 4

OpenStudy (anonymous):

Excellent. This system of equations has two solutions (-1, 1) and (2, 4). Both these solutions satisfy both equations. Well done.

OpenStudy (anonymous):

thank yuoi

OpenStudy (anonymous):

You're welcome.